OpenStudy (anonymous):

I feel really dumb, but am having a brain fart. Simplify, 12-1/9 divided by 4/3 + (1/2)^2

OpenStudy (across):

This?\[\frac{12-\frac{1}{9}}{\frac{4}{3}+\left(\frac{1}{2}\right)^2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

stop cursing yourself or thinking stupid about yourself

OpenStudy (across):

1. Simplify the exponent. 2. Equate bases. 3. Divide.

OpenStudy (anonymous):

well the division sign is between the 1/9 and the 4/3 will it make the problem different if we solved it the way across has it?

OpenStudy (anonymous):